Please find X and Y: 2X + 3Y = 16, 5X - 3Y = 44.
Answer by checkley75(3666)

2x+3y=16
5x-3y=44 now add these two equations.
---------------------------------------
7x=60
x=60/7 answer. now substitute (60/7) for x and solve for y
2*60/7+3y=16
120/7+3y=16
3y=16-120/7
3y=(7*16-120)/7
3y=(112-120)/7
3y=-8/7
y=-8/7*1/3
y=-8/21 answer.
proof.
5*60/7-3*-8/21=44
300/7+24/21=44
(300*3+24)/21=44
(900+24)/21=44
924/21=44
44=44